# Towards Better Comprehension of Breaking Changes in the NPM Ecosystem

### v1

@@ 3.1. Data Collection
Due to the vast number of JavaScript projects in NPM, our study focuses on selecting the most popular projects to ensure that the collected BCs are representative. To do so, we utilize the Libraries.io open source repository and dependency metadata provided by Reid et al. (Reid, 2020) to retrieve the most popular JavaScript and TypeScript projects since Libraries.io is commonly used in prior studies (He et al., 2021; Decan et al., 2019, 2018). We first select popular projects having more than 50 GitHub stars and associated with an existent repository (deprecated and unmaintained repositories are not considered), which yields 35,786 repositories. We then randomly select 381 repositories (95% confidence level and 5% margin of error) for further investigation.
Since we consider the BCs that are explicitly confirmed by developers, we try to extract BCs from developers’ documentation, typically commit messages, changelogs, issues and pull request text on GitHub. We notice that Conventional Commits (Con, [n. d.]) provides a standard for writing readable commit messages. According to the standards, if a commit message contains a BREAKING CHANGE section, the commit is identified as a breaking change. For example, the text below shows a typical commit message indicating a breaking change:
@@
We first check whether Conventional Commits are widely used, analyzing the commits from the 381 projects we used and find that We find that 360 out of 381 repositories contain commits that conform to Conventional Commits and 195 projects have over 80% commits that follow Conventional Commits. We use the regular expression below to check Conventional Commit compliance:
@@
This indicates that Conventional Commit is widely used and we can use it as a simple way to extract plenty of BCs. Considering that many projects do not follow this standard, we also try our best to extract BC commits from changelogs of each project (if developers have declared), issues and pull requests on GitHub. For example, the project tj/commander.js does not use Conventional Commits, then we consult the mentioned pull request IDs (and commits related to the PR) in CHANGELOG.md file [^5] https://github.com/tj/commander.js/blob/master/CHANGELOG.md, shown in Figure 1. In this way, we obtain 5,242 commits.
Figure 1. Screenshot of the Changelog of tj/commander.js (Breaking Changes Are Highlighted)
Then two of the authors manually inspect them and remove some of them that are not related to JavaScript production code (e.g., only modifying package.json, Markdown documentation, build scripts, test cases) or contain very long commit messages (over 10 lines) that are difficult to understand. For example, if a commit only states that “drop support for Node.js 14” and modifies the corresponding item in package.json, it will be discarded. Additionally, we do not include the commits that contain multiple breaking change declarations. Specifically, we remove the 965 commits that only updated dependency versions, 456 commits that only changed minimum versions of supported platforms, and 244 commits that only changed build scripts or JSON configuration files. We also remove 865 commits that contain over 10 line commit messages or multiple BREAKING CHANGE declarations. Finally, we retain 2,724 candidate commits.
@@ 3.2. Regression Testing
We use regression testing to detect BCs in our selected projects (381 in total) to check:
@@
    what percentage of detected BCs are well documented,
@@
    what percentage of documented BCs can be detected by regression testing.
To check ①, due to the number of the total commits is too large (1,455,003), we first sample 16,453 commits for further analysis (99% confidence level and 1% margin of error). We then remove the commits in these projects that only contain non-JavaScript code changes (e.g., test code changes, Markdown and HTML documentation changes, dependency updates in package.json). Then for each commit $c$ and its prior commit $c^{\prime}$ we do regression testing following the steps below:
@@
We use the test cases written by developers rather than dependents since 1) we have found that all of the 381 projects provide test suites and developers’ test cases can better express the intended usage of APIs (Dinella et al., 2022), while many of the 381 projects do not have plenty of dependents, 2) some projects are not intended for pragmatic use, especially CLI tools in our projects such as npm/cli, therefore no clients access the APIs in these projects.
To check ②, we apply the steps described above to the 2,724 BC commits from Section 3.1.
@@ 3.3.1. Breaking Change Selection and Categorization
We select feasible BCs for further research from the 2,724 candidate BC commits and categorize them into types described in Section 2.2. Since we need to understand what a commit changed and why developers made it, we require the commit message and other documentation of a commit (including issues, pull requests and changelogs) to contain some explanation of why developers made that commit. We here reuse Tian et al.’s taxonomy of why information in commit messages (Tian et al., 2022). Typical why information include:
@@
    Requirements Illustration: about the usage need, out-of-date statement, etc.
@@ 3.3.2. Breaking Change Distribution
Figure 2 presents the percentages of different types of BCs. It can be obviously seen that Change Behavior BCs make up the largest proportion of all BC types. By contrast, Move BCs make up the smallest percentage. In our further investigation, we mainly focus on Change Behavior BCs.
@@
The projects that contain BCs in our study serve various types of functionalities, including command-line interface utilities, REST API SDKs and Web frontend frameworks. They can be grouped into six categories (shown in Table 2). For example, pnpm/pnpm [^7] https://github.com/pnpm/pnpm is a alternative CLI tool for official NPM implementation, renovatebot/renovate provides automatic project build and aws/aws-cdk [^8] https://github.com/aws/aws-cdk wraps the AWS cloud APIs, gajus/eslint-plugin-flowtype [^9] https://github.com/gajus/eslint-plugin-flowtype is a ESLint plugin, and async/async [^10] https://github.com/async/async is a fundamental utility of asynchronous programming. Table 3 shows the number of involved projects of each type of BCs.
@@ 3.3.3. Breaking Change Labeling
For the 1,519 remaining commits, two of the authors conduct two thematic analyses to analyze ① source code features (for RQ2 and RQ3) and ② reasons behind the commit (for RQ4). We use the guideline recommended by Cruzes et al. (Cruzes and Dyba, 2011). The guideline for the two analyses is shown below. Two of the authors independently perform the steps above, and they hold a series of meetings to solve the agreements.
@@
    We read the related documentation, especially commit messages carefully to understand what the developers want to express.
@@
    We read the documentation again and generate phrases as initial codes that describe the source code features, how developers make breaking changes and the reasons behind the commit.
@@
    We aggregate the codes with similar meanings and generate a theme name to describe each cluster.
@@
    We then review all themes and try to merge the themes with similar semantics, or make the similar themes become sub-items of a new theme.
After we obtain all themes, we assign possible themes to each BC. Note that for simplicity, we assign each retained BC with one reason, which is consistent with the previous works on motivations behind breaking changes (Brito et al., 2020). In this process, we held several meetings to resolve the disagreements on the reason behind each commit since some BC commits might have multiple potential reasons. For example, in the commit 2713380 of reactivex/rxjs [^11] https://github.com/reactivex/rxjs/commit/2713380, developers renamed inspect to audit, the reasons behind this rename operation was presented in Issue #1505121212https://github.com/reactivex/rxjs/issues/1505 and #1387131313https://github.com/reactivex/rxjs/issues/1387. Developers mentioned two possible reasons, i.e., name collision (since inspect conflicts with util.inspect provided by Node.js [^14] See explanation in Issue 1387 of Rxjs.) and synonym replacement (since audit is synonymous to inspect and it is less confusing). By carefully reading and understanding the related issues, two of the authors reach an agreement and assign the reason “to avoid name conflict” to this BC, because if inspect were not used in Node.js native objects, the identifier inspect is still acceptable and does not need to be renamed. In our dataset, the authors have very few disagreements due to multiple potential reasons.
@@ 4.1. RQ1: To What Extent do Detected Breaking Changes and Documented Breaking Changes Overlap?
From 16,453 commits used in regression testing in Section 3.2, we detect 295 BC commits (1.8%), and 276 of them (93.6%) are documented by developers. And for 2,724 BC commits, 2,206 (81%) can be detected by regression testing.
The results illustrate that most detected BCs are well documented, while a proportion (19%) of documented BCs cannot be detected via regression testing since developers might forget to write test cases or current test cases are too general to cover the code modifications.
The undocumented BC commits are possibly due to ① the commit is made too long ago when Conventional Commits was not proposed, and ② the project does not comply to Conventional Commits. And the main reasons that regression testing cannot detect BCs is that the BC behavior can only be triggered by external conditions (e.g., network error), hence it is difficult to simulate such a situation. For example, in commit 61e7a81a of octokit/rest.js, the “followRedirects” option is no longer supported. This BC can only manifest itself when the HTTP server returns a status code between 301 and 307. However, this is not often seen, and developers have not written any test case to cover the process of “followRedirects” option. Developers might forget to write test cases and current test cases might be too general to cover the code modifications, hence automatic test case generation may be useful.
For the 19% of BCs that cannot be detected, since developers in popular projects have rich development experience, we believe those BCs are trustable and worth attention. Although some BCs currently do not affect many downstream projects, they should still be pointed out since in some cases downstream developers may write code that invokes the broken API, then the downstream developers can confirm the breaking changes quickly by looking up commit messages. For example, in commit 669592d of Socket.io [^15] https://github.com/socketio/socket.io/commit/669592d contains a breaking change that removed Socket#binary method. It is not detected when we run regression testing. However, after several months, one developer asked for an alternative method of Socket#binary [^16] https://github.com/socketio/socket.io/discussions/3826. This indicates that although an API in a BC is not frequently used, and currently no test cases can trigger it, the API is possibly invoked at a certain time.
@@ 4.2. RQ2: What Syntactic Breaking Changes in NPM Ecosystem Are Specific to JavaScript and TypeScript?
While most code-level actions in BCs are also available in other programming languages, such as changing parameter orders and removing classes, etc., we highlight the noticeable syntactic BC actions in our investigated projects.
@@ 4.2.1. JavaScript-specific remove operations
In ECMAScript 6, a module can have many exported items and a default export item. Therefore, developers can not only remove the code of a class, interface or enum, etc. like other OOP languages, but also just remove them from exported list and put them into internal source code files. We present the JavaScript-specific remove operations in our collected BC commits as follows:
  1. (1)
    Remove a default export (6/211 cases). For example, in commit 102e4b0 of nodkz/mongodb-memory-server, developers simply removed default export of the module util. Therefore, the code import generateDbName from 'util' will not work.
  2. (2)
    Remove an export position (17/211 cases). For example in commit cfbfaac of reactivex/rxjs, the types in rxjs/interfaces modules are no longer accessible and users must use them by importing them from the main module. Therefore, users can import UnaryFunction with the following two methods, while after the code change, they can only import UnaryFunction using the second form.
    import { UnaryFunction } from ’rxjs/interfaces’;
    import { UnaryFunction } from ’rxjs’;
@@ 4.2.2. JavaScript-specific signature changes
  1. (1)
    About parameters in configuration objects (15/134 cases). In JavaScript, many properties can be packaged in one configuration object. Considering the diff shown below, the second parameter of the method “userinfo” is an object, and this function only extracts four properties in this object using destructing assignment syntax (des, [n. d.]). After the code change, the “verb” property is renamed to “method”, since “GET” is an HTTP method. Correspondingly, the internal code related to this property is also changed.
@@
    - verb=’GET’, via=’header’, tokenType, params,
    + method=’GET’, via=’header’, tokenType, params,
  2. (2)
    About this parameter (6/134 cases). In the old implementation, the method accepts a callable object (typically a function) and this, and binds the function to this171717bind is a native function in JavaScript. See https://developer.mozilla.org/en-US/docs/Web/JavaScript/Reference/Global_Objects/Function/bind. After the code change, the method no longer provide this parameter, the user should first bind the callable object to this. For example, in the code snippet shown below, findIndex only accepted predicate parameter.
@@
    - return operate(createFind(predicate, thisArg, ’index’));
@@
  3. (3)
    About sync and async (20/134 cases). In JavaScript, asynchronous functions can be implemented by using callback arguments and Promise objects. Legacy JavaScript asynchronous functions are designed with callback-style interface, while new implementations often use async directly. For example, in the commit 8c3cecae of Mongoose, developers provided multiple signatures of an API before the code change, e.g., createIndex have the following function signature:
@@
    - SomeFunction(param, (returnValue, err) => {
    - // Some code about returnValue
@@
  4. (4)
    About undefined and null (8/134 cases). The keyword null represents an intentional empty value, while undefined occurs in accessing uninitialized variables or non-existent object properties, and should not be explicitly assigned to an object. The diff below shows this pattern:
    - public get(key: Key): Value | null {
    - return this.collection.get(key) || null;
@@
  5. (5)
    Switch between “required” and “optional” of fields in method or interface declarations (3/134 cases). TypeScript script projects and part of JavaScript projects use d.ts files to declare interface and method signature. For example, if one parameter is changed to be required, then a “?” will be removed from the parameter type. The diff below (from commit 86074a6 of coinbase/rest-hooks) shows the signature change type:
@@
    - urlParams?: Partial<AbstractInstanceType<T>>,
@@
    - if (urlParams) {

### v2

@@ 3.1. Data Collection
Due to the vast number of JavaScript projects in NPM, our study focuses on selecting the most popular projects to ensure that the collected breaking changes are representative. To do so, we utilize the Libraries.io open source repository and dependency metadata provided by Reid et al. (Reid, 2020) to retrieve the most popular JavaScript and TypeScript projects since Libraries.io is commonly used in prior studies (He et al., 2021; Decan et al., 2019, 2018). We first select popular projects having more than 50 GitHub stars and associated with an existent GitHub repository, which yields 35,786 repositories. We then randomly select 381 repositories (95% confidence level and 5% margin of error) for further investigation.
Since we consider the breaking changes that are explicitly confirmed by developers, we try to extract breaking changes from developers’ documentation, typically commit messages, changelogs, issues and pull request text on GitHub. We notice that Conventional Commits (Con, [n. d.]) provides a standard for writing readable commit messages. According to the standards, if a commit message contains a BREAKING CHANGE section, the commit is identified as a breaking change. For example, the text below shows a typical commit message indicating a breaking change:
@@
We first check whether Conventional Commits are widely used, analyzing the commits from the 381 projects we used. We find that 360 out of 381 repositories contain commits that conform to Conventional Commits and 198 projects have over 80% commits that follow Conventional Commits. We use the regular expression below to check Conventional Commit compliance:
@@
Here fix, feat, chore, etc., are the defined or recommended scope tokens in Conventional Commits specification. This indicates that Conventional Commits specification is widely used and we can use it as a simple way to extract plenty of breaking changes. Considering that many projects do not follow this standard, we also try our best to extract BC commits from changelogs of each project (if developers have declared), issues and pull requests on GitHub. A changelog summarizes the changes from the last version release, and may mention multiple breaking changes. By analyzing the changelog, we can link the mentioned breaking changes to corresponding commits. For example, the project tj/commander.js does not use Conventional Commits, then we consult the mentioned pull request IDs in CHANGELOG.md file [^5] https://github.com/tj/commander.js/blob/master/CHANGELOG.md (shown in Figure 1, where four breaking changes are mentioned), and we also analyze the content of the pull requests to obtain the commits related to the breaking changes. In this way, we obtain 5,242 commits.
Figure 1. Screenshot of the Changelog of commander.js 8.0.0 (The four highlighted breaking changes can be linked to related issues and commits)
Then two of the authors manually inspect them and remove some of them that are not related to JavaScript production code (e.g., only modifying package.json, Markdown documentation, build scripts, test cases) or contain very long commit messages (over 10 lines) that are difficult to understand. For example, if a commit only states that “drop support for Node.js 14” and modifies the corresponding item in package.json, it will be discarded. Additionally, we do not include the commits that contain multiple breaking change declarations. Specifically, we remove 1,826 commits that contain over 10 line commit messages or multiple BREAKING CHANGE declarations and 692 commits that are not JavaScript code changes. Finally, we retain 2,724 candidate commits.
@@ 3.2. Regression Testing
We use regression testing to detect breaking changes in our selected projects (381 in total) to check:
@@
    what percentage of detected breaking changes are well documented,
@@
    what percentage of documented breaking changes can be detected by regression testing.
To check ①, due to the number of the total commits being too large (1,005,344), we first sample 16,371 commits for further analysis (99% confidence level and 1% margin of error). We then remove the commits in these projects that only contain non-JavaScript code changes (e.g., test code changes, Markdown and HTML documentation changes, dependency updates in package.json). Then for each commit $c$ and its prior commit $c^{\prime}$ we do regression testing following the steps below:
@@
We use the test cases written by developers rather than dependents since 1) we have found that 363 out of the 381 projects provide test suites, and developers’ test cases can better express the intended usage of APIs (Dinella et al., 2022), while many of the 381 projects do not have plenty of dependents, 2) some projects are not intended for pragmatic use, especially CLI tools in our projects such as npm/cli, therefore no clients access the APIs in these projects. To check ②, we also apply the steps described above to the 2,724 BC-related commits from Section 3.1.
@@ 3.3.1. Breaking Change Selection and Categorization
We select feasible breaking changes for further research from the 2,724 candidate BC commits and categorize them into types described in Section 2.2. Since we need to understand what a commit changed and why developers made it, we require the commit message and other documentation of a commit (including issues, pull requests and changelogs) to contain some explanation of why developers made that commit. We here reuse Tian et al.’s taxonomy of why information in commit messages (Tian et al., 2022). Typical why information includes:
@@
    Requirements Illustration: about the usage need, out-of-date statements, etc.
@@ 3.3.2. Breaking Change Distribution
Figure 2 presents the percentages of different types of breaking changes. It can be obviously seen that Change Behavior BCs make up the largest proportion of all BC types (68.1%). By contrast, Move and Inline BCs make up the smallest percentage (1.8% and 0.1% respectively). In our further investigation, we mainly focus on Change Behavior BCs.
@@
The projects that contain breaking changes in our study serve various types of functionalities, including command-line interface utilities, REST API SDKs and Web frontend frameworks. They can be grouped into six categories (shown in Table 2). For example, pnpm/pnpm [^7] https://github.com/pnpm/pnpm is a alternative CLI tool for official NPM implementation, renovatebot/renovate provides automatic project build and aws/aws-cdk [^8] https://github.com/aws/aws-cdk wraps the AWS cloud APIs, gajus/eslint-plugin-flowtype [^9] https://github.com/gajus/eslint-plugin-flowtype is an ESLint plugin, and async/async [^10] https://github.com/async/async is a fundamental utility of asynchronous programming. Table 3 shows the number of involved projects for each type of breaking changes.
@@ 3.3.3. Breaking Change Labeling
For the 1,519 remaining commits, two of the authors conducted two thematic analyses to analyze ① source code features (for RQ2 and RQ3) and ② reasons behind the commit (for RQ4). The guideline recommended by Cruzes et al. (Cruzes and Dyba, 2011) were used. The guideline for the two analyses is shown below. Two of the authors independently performed the steps above, and they held a series of meetings to solve the agreements.
@@
    The authors read the related documentation, especially commit messages carefully to understand what the developers want to express.
@@
    The authors read the documentation again and generate phrases as initial codes that describe the source code features, how developers make breaking changes and the reasons behind the commit.
@@
    The authors aggregated the codes with similar meanings and generate a theme name to describe each cluster.
@@
    The authors then reviewed all themes and try to merge the themes with similar semantics, or made the similar themes become sub-items of a new theme.
After obtaining all themes, the authors assigned possible themes to each breaking change. Note that for simplicity, we assigned each retained breaking change with one reason, which is consistent with the previous works on motivations behind breaking changes (Brito et al., 2020). In this process, the authors held several meetings to resolve the disagreements on the reason behind each commit since the reasons behind some BC commits are difficult to determine. For example, in the commit 2713380 of reactivex/rxjs [^11] https://github.com/reactivex/rxjs/commit/2713380, developers renamed inspect to audit, the reasons behind this rename operation was presented in Issue #1505121212https://github.com/reactivex/rxjs/issues/1505 and #1387131313https://github.com/reactivex/rxjs/issues/1387. Developers mentioned two possible reasons, i.e., name collision (since inspect conflicts with util.inspect provided by Node.js [^14] See explanation in Issue 1387 of Rxjs.) and synonym replacement (since audit is synonymous to inspect and it is less confusing). By carefully reading and understanding the related issues, two of the authors reach an agreement and assign the reason “to avoid name conflict” to this breaking change, because if inspect were not used in Node.js native objects, the identifier inspect is still acceptable and does not need to be renamed. In the process of assigning the reason to the breaking changes, the authors have very few disagreements (the Cohen’s Kappa value 0.94).
@@ 4.1. RQ1: To What Extent do Detected Breaking Changes and Documented Breaking Changes Overlap?
From 16,371 commits used in regression testing in Section 3.2, we detect 173 BC commits (1.0%), 95.4% of them (165) are documented by developers. And for 2,724 BC commits, 2,206 (81%) can be detected by regression testing. The results illustrate that most detected breaking changes are well documented, while a proportion (19%) of documented breaking changes cannot be detected via regression testing since developers might forget to write test cases or current test cases are too general to cover the code modifications.
The undocumented breaking change commits are possibly due to ① the commit is made too long ago when Conventional Commits was not proposed, and ② the project does not comply to Conventional Commits. And the main reasons that regression testing cannot detect breaking changes is that the BC behavior can only be triggered by external conditions (e.g., network error), hence it is difficult to simulate such a situation. For example, in commit 61e7a81a of octokit/rest.js, the followRedirects option is no longer supported. This BC can only manifest itself when the HTTP server returns a status code between 301 and 307. However, this is not often seen, and developers have not written any test case to cover the process of “followRedirects” option. Developers might forget to write test cases and current test cases might be too general to cover the code modifications, hence automatic test case generation may be useful.
For the 19% of breaking changes that cannot be detected, since developers in popular projects have rich development experience, we believe those breaking changes are trustable and worth attention. Although some breaking changes currently do not affect many downstream projects, they should still be pointed out since in some cases downstream developers may write code that invokes the broken API, then the downstream developers can confirm the breaking changes quickly by looking up commit messages. For example, in commit 669592d of socketio/socket.io [^15] https://github.com/socketio/socket.io/commit/669592d contains a breaking change that removed Socket#binary method. It is not detected when we run regression testing. However, after several months, one developer asked for an alternative method of Socket#binary [^16] https://github.com/socketio/socket.io/discussions/3826. This indicates that although an API in a breaking change is not frequently used, and currently no test cases can trigger it, the API is possibly invoked at a certain time.
@@ 4.2. RQ2: What Syntactic Breaking Changes in NPM Ecosystem Are Specific to JavaScript and TypeScript?
While most code-level actions in breaking changes are also available in other programming languages, such as changing parameter orders and removing classes, etc., we highlight the noticeable syntactic BC actions in our investigated projects.
@@ 4.2.1. JavaScript-specific remove operations
In ECMAScript 6, a module can have many exported items and a default export item. Therefore, developers can not only remove the code of a class, interface or enum, etc. like other OOP languages, but also just remove them from exported list and put them into internal source code files. We present the JavaScript-specific Remove operations in our collected breaking changes as follows:
[…]
Remove a default export (6 cases). ECMAScript 6 supports default export functionality [^17] https://developer.mozilla.org/en-US/docs/Web/JavaScript/Reference/Statements/export. For example, in commit 102e4b0 of nodkz/mongodb-memory-server, developers simply removed default export of the module util. Therefore, the code import generateDbName from 'util' will not work.
[…]
Remove an export position (17 cases). For example, in commit cfbfaac of reactivex/rxjs, the types in rxjs/interfaces modules are no longer accessible and users must use them by importing them from the main module. Therefore, users can import UnaryFunction with the following two methods, while after the code change, they can only import UnaryFunction using the second form.
1import { UnaryFunction } from ’rxjs/interfaces’;
2import { UnaryFunction } from ’rxjs’;
@@ 4.2.2. JavaScript-specific signature changes
[…]
About parameters in configuration objects (15 cases). In JavaScript, many properties can be packaged in one configuration object. Considering the diff shown below, the second parameter of the method userinfo is an object, and this function only extracts four properties in this object using destructuring assignment syntax (des, [n. d.]). After the code change, the verb property is renamed to method, since GET is an HTTP method. Correspondingly, the internal code related to this property is also changed.
@@
- { verb=’GET’, via=’header’, tokenType, params } = {}
+ { method=’GET’, via=’header’, tokenType, params } = {}
[…]
About this parameter (6 cases). In the old implementation, the method accepts a callable object (typically a function) and this, and binds the function to this181818bind is a native function in JavaScript. See https://developer.mozilla.org/en-US/docs/Web/JavaScript/Reference/Global_Objects/Function/bind. After the code change, the method no longer provide this parameter, the user should first bind the callable object to this. For example, in the code snippet shown below, findIndex only accepted predicate parameter.
@@
- return operate(createFind(predicate, thisArg, ’index’));
@@
[…]
About sync and async (20 cases). In JavaScript, asynchronous functions can be implemented by using callback arguments and Promise objects. Legacy JavaScript asynchronous functions are designed with callback-style interface, while new implementations often use async directly. For example, in the commit 8c3cecae of automattic/mongoose, developers provided multiple signatures of an API before the code change, e.g., createIndex have the following function signature:
@@
- SomeFunction(param, (returnValue, err) => {
- // Some code about returnValue
@@
[…]
About undefined and null (8 cases). The keyword null represents an intentional empty value, while undefined occurs in accessing uninitialized variables or non-existent object properties, and should not be explicitly assigned to an object. The diff below shows this pattern:
- public get(key: Key): Value | null {
- return this.collection.get(key) || null;
@@
[…]
Switch between “required” and “optional” of fields in method or interface declarations (3 cases). TypeScript script projects and part of JavaScript projects use d.ts files to declare interface and method signature. For example, if one parameter is changed to be required, then a “?” will be removed from the parameter type. The diff below (from commit 86074a6 of coinbase/rest-hooks) shows the signature change type:
@@
- urlParams?: Partial<AbstractInstanceType<T>>,
@@
- if (urlParams) {
- if (
- Object.prototype.hasOwnProperty.call(urlParams, ’url’) &&